OpenStudy (anonymous):

f(x) = 4|x + 1| - 6 What is the vertex of the graph of f(x)? (-1, -6) (1, -6) (4, -6)<==My answer

jimthompson5910 (jim_thompson5910):

hint: in general y = a|x-h| + k has the vertex (h,k)

jimthompson5910 (jim_thompson5910):

and it may help to think of x+1 as x - (-1)

OpenStudy (anonymous):

So (4,-6) is right?

jimthompson5910 (jim_thompson5910):

f(x) = 4|x + 1| - 6 y = 4|x + 1| - 6 y = 4|x - (-1)| - 6 What are 'a', h and k?

OpenStudy (anonymous):

a= 4, h= 1, -(-1), k=-6

jimthompson5910 (jim_thompson5910):

h = -1 actually

jimthompson5910 (jim_thompson5910):

if we had |x-1|, then h = 1 but we have |x-(-1)| instead, so that's why h = -1

jimthompson5910 (jim_thompson5910):

The vertex is (h,k), so because h = -1, k = -6, the vertex is (-1,-6)

OpenStudy (anonymous):

okay that makes sense! THanks!

jimthompson5910 (jim_thompson5910):

I'm glad it does, yw